OpenStudy (anonymous):

What is the simplest form of the quotient? ^4sqrt810/^4sqrt2 1)^4sqrt162 2)3^4sqrt5 3)5^4sqrt3 4)3sqrt5

OpenStudy (anonymous):

can you write it down

OpenStudy (anonymous):

i have written down???

OpenStudy (anonymous):

\[ \frac{\sqrt[4]{810}}{\sqrt[4]{2}} \]Step one: use \(\color{blue}a^c / \color{red}b^c = (\color{blue}a /\color{red}b)^c\)\[ \frac{\sqrt[4]{\color{blue}{810}}}{\sqrt[4]{\color{red}2}} = \sqrt[4]{\frac{\color{blue}{810}}{\color{red}2}} \]Step two: \[ \sqrt[4]{\frac{810}{2}}=\sqrt[4]{405} \]Step three: We need to factor. Can you factor?

OpenStudy (anonymous):

Facotr what? ^4sqrt405?

OpenStudy (anonymous):

Factor 405 into prime numbers.

OpenStudy (anonymous):

would it be factored by 5?

OpenStudy (anonymous):

yeah that is one factor... keep going.

OpenStudy (anonymous):

10?

OpenStudy (anonymous):

\[ 405 = 81 \times 5= 27\times 3\times 5=9\times 3\times 3\times 5=3\times 3\times 3\times 3\times 5 = 3^4\times 5 \]So

OpenStudy (anonymous):

So that means \[ \sqrt[4]{405} =\sqrt[4]{ 3^4\times 5} = 3\sqrt[4]{5} \]